# Robust Amortized Bayesian Inference with Self-Consistency Losses on Unlabeled Data

### v1

@@ 2.3 Self-Consistency Losses Are Strictly Proper
The strict properness extends to semi-supervised losses of the form (2), which combine standard simulation-based losses with self-consistency losses.
Proposition 3.
Proposition 4.
$C\left(\frac{q(x\mid\theta)\,p(\theta)}{p(\theta\mid x)}\right)$ (12)
are strictly proper as well. That is, they are globally minimized if and only if $q(x\mid\theta)=p(x\mid\theta)$ almost everywhere.
The proof of Proposition 4 proceeds in the same manner as for Proposition 1, just exchanging likelihood and posterior.
Clearly, strict properness does not necessarily hold if both posterior and likelihood are unknown or approximate. This is because any pair of approximators $q(\theta\mid x)$ and $q(x\mid\theta)$ that satisfy $q(\theta\mid x)\propto q(x\mid\theta)\,p(\theta)$ minimize the self-consistency loss regardless of their relation to the accurate posterior $p(\theta\mid x)$ and likelihood $p(x\mid\theta)$. For example, the choices $q(\theta\mid x)=p(\theta)$ and $q(x\mid\theta)\propto 1$ minimize the self-consistency loss, but may be arbitrarily far away from their actual target distributions.
In other words, if both likelihood and posterior are unknown, the self-consistency loss needs to be coupled with another loss component, such as the maximum likelihood loss, to enable joint learning of both approximators $q(\theta\mid x)$ and $q(x\mid\theta)$ [34]. Nevertheless, the self-consistency loss still yields significant improvements: as demonstrated in our experiments, the semi-supervised loss (2) significantly enhances the robustness of ABI even when both the posterior and likelihood are unknown.
3 Related Work
The robustness of ABI and simulation-based inference methods more generally has been the focus of multiple recent studies [e.g., 11, 10, 5, 37, 16, 21, 12, 35, 23, 38, 30, 33]. These efforts can be broadly classified into two categories: (a) analyzing or detecting simulation gaps and (b) mitigating the impact of simulation gaps on posterior estimates. Since our work falls into the latter category, we briefly discuss methods aimed at increasing the robustness of fully amortized approaches. For instance, Gloeckler et al. [16] explore efficient regularization techniques that trade off some posterior accuracy to enhance the robustness of posterior estimators against adversarial attacks. Ward et al. [37] and Siahkoohi et al. [35] apply post hoc corrections based on real data, utilizing MCMC and the reverse Kullback-Leibler divergence, respectively. Differently, Gao et al. [12] propose a departure from standard Bayesian inference by minimizing the expected distance between simulations and observed data, akin to generalized Bayesian inference with scoring rules [30]. Perhaps the closest work in spirit to ours is Wehenkel et al. [38], which introduces the use of additional training information in the form of a (labeled) calibration set $(x^{*},\theta^{*})$ which contains observables from the real data distribution as well as the corresponding ground-truth parameters.
Figure 2: Bias of posterior mean, bias of posterior standard deviation [13] and posterior distance quantified by maximum mean discrepancy [19] to the analytic posterior for variations of the default configuration outlined in Section 4. NPE approximators with the added self-consistency loss component are shown in blue, NPE approximators using just the standard simulation-based loss are shown in red. Irrespective of the varied factor and for all metrics, adding the self-consistency loss component is always a drastic improvement over the standard simulation-based loss alone. The plots show that adding the self-consistency loss component provides strong robustness gains even in high-dimensional spaces (top row) or when the self-consistency loss is evaluated on little data (center row). Variation of the mean of the unlabeled training data show that adding the self-consistency loss drastically improves posterior estimation as long as data used for evaluating the self-consistency loss is at least slightly out-of-distribution compared to the original training data ($\mu^{*}\geq 1$).
4 Case Study: Multivariate Gaussian Models
We illustrate the usefulness of our proposed self-consistency loss on the following multivariate Gaussian model, also known as the normal means problem. This model has also been considered in previous studies investigating the robustness of neural ABI [33]. It allows us to explore the behavior of NPE trained with the semi-supervised loss (2) in a simple, controlled setting where the true posterior is available in closed-form. The prior and likelihood are given by
[…]
The parameters $\theta\in\mathbb{R}^{D}$ are sampled from a $D$-dimensional multivariate Gaussian distribution with mean vector $\mu_{\rm prior}$ and diagonal covariance matrix $\sigma^{2}_{\rm prior}\,I_{D}$. Here, we fix $\mu_{\rm prior}=0$ and $\sigma^{2}_{\rm prior}=1$. On this basis, $K$ independent, synthetic data points $x^{(k)}\in\mathbb{R}^{D}$ are sampled from a $D$-dimensional multivariate Gaussian distribution with mean vector $\theta$ and diagonal covariance matrix $\sigma^{2}_{\rm lik}\,I_{D}$. We fix $\sigma^{2}_{\rm lik}=K$ such that the total information in $x$ remains constant, independent of $K$, which simplifies comparisons across observations of varying number of data points.
@@ 4 Case Study: Multivariate Gaussian Models
From this model, we simulate a labeled training dataset with a budget of $N=1024$, that is, $N$ independent instances of $\theta_{n}$ (the "labels") with corresponding observations $x_{n}=\{x_{n}^{(k)}\}_{k=1}^{K}$, each consisting of $K$ data points. This labeled training dataset $\{(\theta_{n},x_{n})\}_{n=1}^{N}$ is used for optimizing the standard simulation-based loss component. The self-consistency loss component is optimized on an additional unlabeled dataset $\{x^{*}_{m}\}_{m=1}^{M}$ of $M=32$ independent sequences $x^{*}_{m}=\{x_{m}^{*(k)}\}_{k=1}^{K}$, which, for the purpose of this case study, are simulated from
@@
The faithfulness of the approximated posteriors $q(\theta\mid x_{\rm obs})$ are assessed by computing the bias in posterior mean and standard deviation as well as the maximum mean discrepancy (MMD) with a Gaussian kernel [19] between the approximate and true (analytic) posterior. Mathematical details of the analytic posterior can be found in Appendix A.
In our numerical experiments, we study the influence of several aspects of the normal model on the performance of NPE. To prevent combinatorial explosion, we vary the factors below separately, with all other factors fixed to their default configuration (highlighted in bold):
  • parameter dimensionality ($D=2,\textbf{10},100$)
  • number of unlabeled observations for the self-consistency loss $\{{x^{*}_{m}}\}_{m=1}^{M}$ ($M=1,4,\textbf{32})$
  • mean $\mu^{*}$ of the unlabeled observations $x^{*}_{m}$ ($\mu^{*}=0,1,2,\textbf{3},5$)
  • inclusion of a summary network $(K=10)$ or not $\bm{(K=1)}$
  • likelihood function (known, estimated)
For the neural posterior estimators $q(\theta\mid x)$, we use a neural spline flow [7] with 5 coupling layers of 128 units each utilizing ReLU activation functions, L2 weight regularization with factor $\gamma=10^{-3}$, 5% dropout and a multivariate unit Gaussian latent space. The network is trained using the Adam optimizer for 100 epochs with a batch size of 32 and a learning rate of $5\times 10^{-4}$. These settings were the same for both the standard simulation-based loss and our proposed semi-supervised loss. For the conditions involving an estimated likelihood $q(x\mid\theta)$, we use the same configuration for the likelihood network as for the posterior network. For the summary network $h(x)$ (if included), we use a deep set architecture [39] with 30 summary dimensions and mean pooling, 2 equivariant layers each consisting of 2 dense layers with 64 units and a ReLU activation function. The inner and outer pooling functions also use 2 dense layers with the same configuration. The likelihood network as well as the summary network are jointly trained with the inference network using the Adam optimizer for 100 epochs with a batch size of 32 and a learning rate of $5\times 10^{-4}$.
4.1 Results
In Figure 1, we depict the results obtained from (a) standard NPE (trained on the simulation-based loss only), (b) our semi-supervised NPE (adding the self-consistency loss with known likelihood) and, (c) the gold-standard (analytic) reference. We see that standard NPE already completely fails for $x_{\rm obs}\sim N(\mu_{\rm obs}=2,0.01I_{D})$, and subsequently also for any larger values $\mu_{\rm obs}>2$. In contrast, adding the self-consistency loss component to obtain our semi-supervised loss achieves almost perfect posterior estimation. This holds true even in cases where $x_{\rm obs}$ is multiple standard deviations away from all the training data, that is, from both the labeled dataset $\{(\theta_{n},x_{n})\}_{n=1}^{N}$ and the unlabeled dataset $\{{x^{*}_{m}}\}_{m=1}^{M}$. These results indicate that the self-consistency criterion can provide strong robustness gains even far outside the typical space of the training data.
In Figure 2, we show posterior mean and standard deviation bias as well as the MMD between the approximate and true posterior for each of the above five factors. When varying the parameter dimensionality, the results indicate that including the self-consistency loss component achieves almost perfect posterior approximation for up to 10 dimensions even in the presence of extreme deviations from the initial training data. While we do see larger discrepancies from the true posterior for 100 dimensions, adding the self-consistency loss still considerably improves estimates. The dataset size factor shows strong robustness gains even when few data are used to evaluate the self-consistency loss (4 unlabeled vs. 1024 labeled training observations). We already see clear improvements over the standard simulation-based loss when only a single unlabeled observation is seen during training. Varying the mean $\mu_{\rm obs}$ of the new observations shows that, as long as the data used for evaluation the self-consistency loss is not identical to the training data (i.e., as long as $\mu_{\rm obs}\neq 0$), including the self-consistency loss component enables accurate posterior approximation far outside the typical space of the training data.
In Figure 3, we see that the benefits of self-consistency persist when the posterior is conditioned on more than one data point per observation (i.e., when $K=10$), that is, in the presence of a summary network. Further, we still see clear benefits of adding the self-consistency loss even when the likelihood is estimated via by a neural likelihood approximator $q(x\mid\theta)$, trained jointly with the neural posterior approximator $q(\theta\mid x)$ on the same training data. However, with an estimated likelihood, posterior bias, especially bias in the posterior standard deviation, and MMD distance to the true posterior are larger than in the known likelihood case.
5 Discussion
In this proof-of-concept paper, we demonstrated that Bayesian self-consistency losses can drastically improve the robustness of neural amortized Bayesian inference (ABI) on out-of-simulation data. Performing accurate inference on such data – outside of the space seen during simulation-based training (e.g., because of model misspecification) – has been one of the major challenges of ABI since its inception: conventional ABI approaches are known to dramatically fail in such cases [33, 16, 21], as we also illustrated in our experiments. In contrast, when adding the self-consistency loss and training it on unlabeled out-of-simulation data, we obtained near-unbiased posterior estimation far beyond the training data spaces. That is, ABI became highly accurate even for data sets far away from any data seen during training by either of the two loss components – the standard simulation-based loss and the self-consistency loss. Importantly, because the self-consistency loss does not require data labels (i.e., knowledge of true parameter values), we can also use any amount of real data during training to improve the robustness of ABI.
We believe that the self-consistency loss trained on out-of-simulation data is a major breakthrough in neural ABI. We had originally proposed this loss to improve training efficiency for models with slow simulators [34, 22], but had not considered it in combination with (potentially real) out-of-simulation data. To the best of our knowledge, no existing ABI approach achieves similar degrees of robustness as our method presented here. The strong robustness gains persisted even in relatively high dimensional models (we tested a model with up to 100 parameters), suggesting no major curse of dimensionality.
Also, when using a learned rather than known likelihood function, robustness was clearly improved. However, the resulting amortized inference with a learned likelihood did not (yet) achieve the same level of accuracy on out-of-simulation data as the known likelihood case. Additionally, our variance-based self-consistency loss relies on fast density evaluations during training to maintain competitive training times. This currently makes free-form methods, such as flow matching [27], less practical due to their reliance on numerical integration for density estimation. As a result, the joint learning of posteriors and likelihoods, along with the development of efficient self-consistency losses for free-form flows, remains an open area for future research.
6 Acknowledgments
Daniel Habermann, Stefan Radev, and Paul Bürkner acknowledge support of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Projects 508399956 and 528702768. Paul Bürkner further acknowledges support of the DFG Collaborative Research Center 391 (Spatio-Temporal Statistics for the Transition of Energy and Transport) – 520388526.
Appendix
Appendix A: Analytic posterior of the normal means problem
The analytic posterior for the for the normal means problem introduced in Section 4 is a conjugate normal distribution

### v2

@@ A.1 Proofs
A.2 Detailed setup of the multivariate normal case study
@@
From the multivariate normal model described in Section 4.1, we simulate a labeled training dataset with a budget of $N=1024$, that is, $N$ independent instances of $\theta_{n}$ (the "labels") with corresponding observations $x_{n}=\{x_{n}^{(k)}\}_{k=1}^{K}$, each consisting of $K$ data points. This labeled training dataset $\{(\theta_{n},x_{n})\}_{n=1}^{N}$ is used for optimizing the standard simulation-based loss component. The self-consistency loss component is optimized on an additional unlabeled dataset $\{x^{*}_{m}\}_{m=1}^{M}$ of $M=32$ independent sequences $x^{*}_{m}=\{x_{m}^{*(k)}\}_{k=1}^{K}$, which, for the purpose of this case study, are simulated from
@@
The faithfulness of the approximated posteriors $q(\theta\mid x_{\rm obs})$ are assessed by computing the bias in posterior mean and standard deviation as well as the maximum mean discrepancy (MMD) with a Gaussian kernel [22] between the approximate and true (analytic) posterior.
The analytic posterior for the normal means problem is a conjugate normal distribution
@@
For the NPEs $q(\theta\mid x)$, we use a neural spline flow [7] with 5 coupling layers of 128 units each utilizing ReLU activation functions, L2 weight regularization with factor $\gamma=10^{-3}$, 5% dropout and a multivariate unit Gaussian latent space. The network is trained using the Adam optimizer for 100 epochs with a batch size of 32 and a learning rate of $5\times 10^{-4}$. These settings were the same for both the standard simulation-based loss and our proposed semi-supervised loss. For the conditions involving an estimated likelihood $q(x\mid\theta)$, we use the same configuration for the likelihood network as for the posterior network. For the summary network $h(x)$ (if included), we use a deep set architecture [43] with 30 summary dimensions and mean pooling, 2 equivariant layers each consisting of 2 dense layers with 64 units and a ReLU activation function. The inner and outer pooling functions also use 2 dense layers with the same configuration. The likelihood network as well as the summary network are jointly trained with the inference network using the Adam optimizer for 100 epochs with a batch size of 32 and a learning rate of $5\times 10^{-4}$.
A.3 Comprehensive results for the multivariate normal case study
Figure 3: Bias of posterior mean, bias of posterior standard deviation and posterior distance quantified by maximum mean discrepancy to the analytic posterior for variations of the default configuration outlined in Section 4.1. NPE approximators with the added self-consistency loss component are shown in blue, NPE approximators using just the standard simulation-based loss are shown in red. Irrespective of the varied factor and for all metrics, adding the self-consistency loss component is always a drastic improvement over the standard simulation-based loss alone. The plots show that adding the self-consistency loss component provides strong robustness gains even in high-dimensional spaces (top row) or when the self-consistency loss is evaluated on little data (center row). Variation of the mean of the unlabeled training data show that adding the self-consistency loss drastically improves posterior estimation as long as data used for evaluating the self-consistency loss is at least slightly out-of-distribution compared to the original training data ($\mu^{*}\geq 1$).
A.4 Detailed setup of the air traffic case study
For the air traffic model defined in Section 4.2, we set independent priors on the parameters as follows:
$\alpha_{j}$ $\sim\mathrm{Normal}(0,0.5)\hskip 42.67912pt\beta_{j}\sim\mathrm{Normal}(0,0.2)$
$\gamma_{j}$ $\sim\mathrm{Normal}(0,0.5)\hskip 42.67912pt\delta_{j}\sim\mathrm{Normal}(0,0.5)$ (20)
$\mathrm{log}(\sigma_{j})$ $\sim\mathrm{Normal}(-1,0.5).$
For the NPEs $q(\theta\mid x)$, we use a neural spline flow [7] with 6 coupling layers of 128 units each utilizing exponential linear unit activation functions, L2 weight regularization with factor $\gamma=10^{-3}$, 5% dropout and a multivariate unit Gaussian latent space. These settings were the same for both the standard simulation-based loss and our proposed semi-supervised loss. The simulation budget was set to $N=1024$. For the summary network, we use a long short-term memory layer with 64 output dimensions followed by two dense layers with output dimensions of 256 and 64. The inference and summary networks are jointly trained using the Adam optimizer for 100 epochs with a batch size of 32 and a learning rate of $5\times 10^{-4}$.
A.5 Comprehensive results for the air traffic case study
Table 2: Posterior metrics for NPE and NPE augmented with self-consistency loss (NPE + SC) relative to Stan. For each parameter, the absolute bias in posterior means and standard deviations are reported along with the Wasserstein distance between the posteriors, using Stan as reference. The self-consistency loss was evaluated on data from $M=$ 4 countries during training. Metrics are averaged over all 15 countries.
Parameter $|\mu-\mu_{\rm Stan}|$ $|\sigma-\sigma_{\rm Stan}|$ Wasserstein distance
NPE NPE+SC NPE NPE+SC NPE NPE+SC
$\alpha$ $0.078$ 0.003 $0.033$ 0.019 $0.085$ 0.032
$\beta$ $0.152$ 0.011 $0.054$ 0.011 $0.161$ 0.070
$\gamma$ $0.087$ 0.047 $0.057$ 0.022 $0.153$ 0.112
$\delta$ $0.052$ 0.046 $0.037$ 0.018 $0.119$ 0.102
$\log(\sigma)$ $0.214$ 0.207 0.049 $0.057$ $0.304$ 0.281
Table 3: Posterior metrics for NPE and NPE augmented with self-consistency loss (NPE + SC) relative to Stan. For each parameter, the absolute bias in posterior means and standard deviations are reported along with the Wasserstein distance between the posteriors, using Stan as reference. The self-consistency loss was evaluated on data from $M=$ 15 countries during training. Metrics are averaged over all 15 countries.
Parameter $|\mu-\mu_{\rm Stan}|$ $|\sigma-\sigma_{\rm Stan}|$ Wasserstein distance
NPE NPE+SC NPE NPE+SC NPE NPE+SC
$\alpha$ $0.078$ 0.002 $0.033$ 0.001 $0.085$ 0.006
$\beta$ $0.152$ 0.001 $0.054$ 0.001 $0.161$ 0.009
$\gamma$ $0.087$ 0.002 $0.057$ 0.006 $0.153$ 0.014
$\delta$ $0.052$ 0.003 $0.037$ 0.005 $0.119$ 0.013
$\log(\sigma)$ $0.214$ 0.002 $0.049$ 0.004 $0.304$ 0.011
Figure 5: Comparison of posterior estimates between standard amortized NPE (red circles), NPE augmented by our self-consistency loss (NPE + SC; blue squares) and Stan (reference; gray triangles). The plots illustrate central 50% (thick lines) and 95% (thin lines) credible intervals of all five parameters for different countries, sorted by the lower 5% quantile according to Stan. Abbreviations follow the ISO 3166 alpha-2 codes. The self-consistency loss was evaluated on data from $M=$ 4 countries during training.
Figure 6: Comparison of posterior estimates between standard amortized NPE (red circles), NPE augmented by our self-consistency loss (NPE + SC; blue squares), and Stan (reference; gray triangles). The plots illustrate central 50% (thick lines) and 95% (thin lines) credible intervals of all five parameters for different countries, sorted by the lower 5% quantile according to Stan. Abbreviations follow the ISO 3166 alpha-2 codes. The self-consistency loss was evaluated on data from $M=$ 8 countries during training.
Figure 7: Comparison of posterior estimates between standard amortized NPE (red circles), NPE augmented by our self-consistency loss (NPE + SC; blue squares), and Stan (reference; gray triangles). The plots illustrate central 50% (thick lines) and 95% (thin lines) credible intervals of all five parameters for different countries, sorted by the lower 5% quantile according to Stan. Abbreviations follow the ISO 3166 alpha-2 codes. The self-consistency loss was evaluated on data from $M=$ 15 countries during training.